OpenStudy (anonymous):

Emily has 250 books. If Frank were to double the number of books that he now owns, he would have more than Emily has.

OpenStudy (anonymous):

That means that Frank has at least 126 books.

OpenStudy (anonymous):

whats the question

OpenStudy (sarahjoy26):

Should there be more information?

OpenStudy (anonymous):

Write a mathematical sentence that expresses the information given below. Use p as your variable name. If necessary: type < = to mean or > = to mean . The original price of a computer was reduced by 250 dollars. The new, discounted price is 1499 dollars. Answer:

OpenStudy (sarahjoy26):

Yeah i agree @Bamawam because 250 divided by 2 is 126 so that means Frank has at least 126 books.

OpenStudy (sarahjoy26):

I mean 125

OpenStudy (sarahjoy26):

Because 125*2 = 250

OpenStudy (anonymous):

wrong because it says when he doubled it he had more than emily

OpenStudy (anonymous):

250/2=125 That means that Frank needs at least 126 books now so that 126*2=254 And as you know 254>250 Anything 254 and above is correct.

OpenStudy (sarahjoy26):

It is saying before he doubled it, how many would he have.

OpenStudy (anonymous):

Sorry, anything 126 and above is correct.

OpenStudy (anonymous):

@Bamawam is correct

OpenStudy (sarahjoy26):

Exactly, @Bamawam because it said MORE than it.

OpenStudy (sarahjoy26):

Wow this is a lot to explain. Ok here goes: Emily has 250 books. Frank has more than her when he doubles his real amount of books. So, his real amount of books would be anything above 125 because 520 divided by 2 (which is reversing the doubled amount of books) is 125. So if he would have more than her if he doubled his books, than before he doubled his books, he would have no less than 125. @Kendraluv2 There you go. :)

OpenStudy (sarahjoy26):

When I said 520 I meant to say 250. LOL :)